A sample of 6 elements was taken and their measurements are below:
39, 22, 31, 25, 33, 30
With 80% confidence, give a formula for a range of values that the population mean is likely to fall in.

Group of answer choices
Answer #1

sample mean, xbar = 30
sample standard deviation, s = 6
sample size, n = 6
degrees of freedom, df = n - 1 = 5

Given CI level is 80%, hence α = 1 - 0.8 = 0.2
α/2 = 0.2/2 = 0.1, tc = t(α/2, df) = 1.476

CI = (x-te x 18 +te x )
CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (30 - 1.476 * 6/sqrt(6) , 30 + 1.476 * 6/sqrt(6))
CI = (26.3846 , 33.6154)

rounding to 2 decimal
26.38 , 33.62